izze2004:

Bodhi has a collection of 175 dimes and nickels. The collection is worth $13.30. Which equation can be used to find n, the number of nickels in the collection? 0.1n + 0.05(n – 175) = 13.30 0.1n + 0.05(175 – n) = 13.30 0.1(n – 175) + 0.05 = 13.30 0.1(175 – n) + 0.05n = 13.30

izze2004:

i really need help and fast

justjm:

denote d=dimes and n=nickels. Make a system of equations. We can say that \(n+d=175\) Recall that n=$0.05 and d=$0.1 We can then say that \(0.05n+0.1d=13.30\) Now you have two systems. It seems like your options have decided to continue using substitution. \((1)...n+d=175\) \((2)...0.05n+0.1d=13.30\) Rearranging equation 1 is much easier than 2, so that's how we can continue. \(d=175-n\) Substitute the expression for d into \(0.05n+0.1d=13.30\) and that should arrive you to your answer.